From Van Buren Bridge I cycle straight east along Hwy34 for 7.2 miles. I turn left into

Looney Lane and ride straight north for 2.0 miles where I reach the LBCC campus.

Along Hwy34, the traffic is sometimes terrifying. A few years ago the new bike path

was built, and I tried it out. I did not take Hwy34 but followed the new safe route that

was suggested by friends. So I rode the 3.1 miles on the new bike path from the

VanBuren/Harrison Bridge along Hwy34 east, then north east along Riverside Dr.,

SW Queen Ave. and finally along Hwy99E to LBCC. It was a Saturday, and I was

30min late to the TASS session at LBCC. That was not what I had planned, and I am

back riding along Hwy34 with my heart fearfully beating but being in time for class.

On my rides I often think about how another bike path could create a better and safer

riding experience. 


a) Assume the community would agree to build a necessary bike path in a straight

line from the intersection of Hwy34 and Riverside Dr. to the LBCC campus.

What would be the distance of that bike ride?


1
Expert's answer
2021-01-27T16:34:33-0500
"d=\sqrt{x^2+y^2}\\d=\sqrt{7.2^2+2.0^2}=7.5\ miles"